# On Creating a Causally Grounded Usable Rating Method for Assessing the Robustness of Foundation Models Supporting Time Series

### v2

@@ Appendix A Details of Rating Algorithms
Appendix B Additional Related Work
B.1 Robustness Testing of Foundation Models
(Zhang and Ré 2022) examined group distribution shifts and evaluated FMs on image classification tasks with spurious confounders. In our work, we assess the robustness of FMs within time series forecasting by measuring their performance in the presence of two confounders across various perturbation settings and test dataset distributions. (Zhang et al. 2023) used foundation models as a surrogate oracle to measure the robustness of image classification models. However, their test systems did not include any foundation models. (Xiao et al. 2024) introduces a framework, RITFIS, to assess the LLM-based software against natural language input. However, they did not consider any other modalities. (Schlarmann and Hein 2023) showed that imperceivable attacks on images to change the caption output of multi-modal FMs can lead to broadcasting of fake information to honest users. They only evaluate robustness of OpenFlamingo model under different attacks but does not compare its performance with any other FMs. None of these works assess the effectiveness and usability of their robustness testing methods. We address this gap through a user study.
Appendix C Evaluation Metrics
In this section, we define our evaluation metrics: forecasting accuracy and robustness.
Forecasting Accuracy Metrics
We evaluate the systems’ forecasting accuracy using established metrics commonly applied in time-series forecasting tasks (Makridakis, Spiliotis, and Assimakopoulos 2022).
Symmetric mean absolute percentage error (SMAPE) is defined as,
$SMAPE=\frac{1}{T}\sum_{t=1}^{T}\frac{|x_{t}-\hat{x}_{t}|}{(|x_{t}|+|\hat{x}_{t}|)/2},$ (3)
$MASE=\frac{\frac{1}{T}\sum_{i=t+1}^{t+T}|x_{i}-\hat{x_{i}}|}{\frac{1}{t}\sum_{i=1}^{t}|x_{i}-x_{i-1}|},$ (4)
Sign Accuracy quantifies the average classification accuracy across all test samples, where a higher accuracy indicates more precise predictions. This metric classifies based on how the predicted forecasts align with the most recent observed values in the input time series.
Robustness Metrics We adapt WRS metric originally proposed in (Lakkaraju, Srivastava, and Valtorta 2024) to answer RQ1. Additionally, we introduce two new metrics: APE and PIE % (modified versions of ATE (Abdia et al. 2017) and DIE % (Lakkaraju, Srivastava, and Valtorta 2024)) tailored to answering RQ2 and RQ3.
Weighted Rejection Score (WRS): WRS, introduced in (Lakkaraju, Srivastava, and Valtorta 2024), measures statistical bias. First, Student’s t-test (Student 1908) compares max residual distributions $(R^{max}_{t}|Z)$ for different values of the protected attribute $Z$. We measure this between each pair of industries or companies, resulting in ${}^{3}C_{2}=3$ comparisons. (Lakkaraju, Srivastava, and Valtorta 2024) chose different confidence intervals (CI) [95%, 75%, 60%] that have different critical values and if the computed t-value is less than the critical value, the null hypothesis is rejected. WRS is mathematically defined by the following equation:
[…]
where, $x_{i}$ is the variable set based on whether the null hypothesis is accepted (0) or rejected (1). $w_{i}$ is the weight that is multiplied by $x_{i}$ based on the CI. For example, if CI is 95%, $x_{1}$ is multiplied by 1. The lower the CI, the lower the weight will be. WRS helps us answer RQ1 (see Section 3).
Appendix D Figures in a Higher Resolution from Main Paper
Figure 7: Causal model $\mathcal{M}$ for FMTS. The validity of link ‘1’ depends on the data distribution ($P|Z$), while the validity of the links ‘2’ and ‘3’ are tested in our experiments.
Figure 8: Variants of the causal diagram in Figure 1 used to answer different research questions (RQs).
Figure 9: Data to predictions. Workflow for performing statistical and causal analysis to compute raw scores and assign final ratings to the test systems
Figure 10: Predictions to ratings. Black arrows denote the unperturbed and red arrows indicate the perturbed paths. Dashed lines shows the multi-modal path. The perturbed parts of the plots are highlighted in red.
Figure 11: Studying each metric with respect to impact of company and industry as confounders for all models and all perturbations. Plotted in double logarithmic scale, lower values indicate better robustness. Ratings generated by our method (with $L=3$) are shown on the top of each plot. The complete final order (with ratings) are shown in Table 2 in Appendix E.
Figure 12: Radar plots showing for all systems (a) mean forecasting accuracy with respect to all metrics, (b) forecasting accuracy under P2 (half-valued pertubation) Appendix E: Figs. 17, 18 show all perturbations; (c) mean robustness metrics for FMTS and $S_{a}$, and (d) robustness under P2. Each axis is normalized and inverted if needed so that outer ring implies better performance.
Figure 13: Effect of the modalities for $S_{g}$ (left) and $S_{p}$ (right).
Figure 14: Role of architecture in forecasting accuracy and robustness. Performance is averaged across models within each category. See Table 1.
Appendix E Additional Experimental Results
In this section, Table 2 shows the partial order and final order with respect to all the metrics defined in Section 4.1. Table 3 shows the forecasting accuracy values for all the systems. Table 4 shows the research questions, and average values for all metrics across systems and perturbations (average values are referred to in the conclusions made for each RQ in Section 5.2). Figure 15 and 16 shows bar plots with all the robustness metric values and forecasting accuracy values. Figure 19 shows the heatmap for all metrics for all the models.
Table 2 shows the partial order and final order with respect to all the metrics defined in Section 4.1. Table 3 shows the forecasting accuracy values for all the systems. Table 4 shows the research questions, and average values for all metrics across systems and perturbations (average values are referred to in the conclusions made for each RQ in Section 5.2). Figure 15 and 16 shows bar plots with all the robustness metric values and forecasting accuracy values. Figure 19 shows the heatmap for all metrics for all the models.
Forecasting Evaluation Dimensions P Partial Order Complete Order
[…]
P1 {$S_{a}$: 2.6, $S_{m}$: 4.6, $S_{g}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{r}$: 4.6, $S_{p}^{ni}$: 5.9, $S_{p}$: 6.9, $S_{c}$: 6.9, $S_{b}$: 6.9} {$S_{a}$: 1, $S_{m}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{r}$: 1, $S_{p}^{ni}$: 2, $S_{p}$: 3, $S_{c}$: 3, $S_{b}$: 3}
P2 {$S_{a}$: 4.6, $S_{g}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{p}^{ni}$: 4.6, $S_{m}$: 4.6, $S_{r}$: 4.6, $S_{c}$: 6.9, $S_{p}$: 6.9, $S_{b}$: 6.9} {$S_{a}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 1, $S_{m}$: 1, $S_{r}$: 1, $S_{c}$: 2, $S_{p}$: 2, $S_{b}$: 2}
[…]
P1 {$S_{a}$: 0.6, $S_{c}$: 4.6, $S_{p}$: 5.9, $S_{p}^{ni}$: 5.9, $S_{r}$: 5.9, $S_{g}^{ni}$: 5.9, $S_{g}$: 6.9, $S_{m}$: 6.9, $S_{b}$: 6.9} {$S_{a}$: 1, $S_{c}$: 1, $S_{p}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 2, $S_{g}^{ni}$: 2, $S_{g}$: 3, $S_{m}$: 3, $S_{b}$: 3}
P2 {$S_{a}$: 2.6, $S_{c}$: 4.6, $S_{r}$: 4.6, $S_{p}^{ni}$: 4.6, $S_{p}$: 5.2, $S_{g}$: 5.9, $S_{g}^{ni}$: 5.9, $S_{m}$: 6.9, $S_{b}$: 6.9} {$S_{a}$: 1, $S_{c}$: 1, $S_{r}$: 1, $S_{p}^{ni}$: 1, $S_{p}$: 2, $S_{g}$: 2, $S_{g}^{ni}$: 2, $S_{m}$: 3, $S_{b}$: 3}
P3 {$S_{g}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{p}^{ni}$: 4.6, Sp: 4.6, $S_{c}$: 4.6, $S_{m}$: 6.9, $S_{a}$: 6.9, $S_{r}$: 6.9, $S_{b}$: 6.9} {$S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 1, Sp: 1, $S_{c}$: 1, $S_{m}$: 2, $S_{a}$: 2, $S_{r}$: 2, $S_{b}$: 2}
[…]
P2 {$S_{c}$: 357.72, $S_{g}^{ni}$: 527.76, $S_{g}$: 597.54, $S_{a}$: 902.54, $S_{m}$: 1326.20, $S_{r}$: 1463.71, $S_{p}^{ni}$: 1653.53, $S_{b}$: 2174.39, $S_{p}$: 2295.68} {$S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{a}$: 2, $S_{m}$: 2, $S_{r}$: 2, $S_{p}^{ni}$: 3, $S_{b}$: 3, $S_{p}$: 3}
P3 {$S_{g}$: 703.94, $S_{g}^{ni}$: 884.34, $S_{c}$: 911.53, $S_{p}^{ni}$: 972.95, $S_{a}$: 1195.04, $S_{m}$: 2998.25, $S_{p}$: 3208.04, $S_{r}$: 3560.94, $S_{b}$: 7489.48} {$S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{c}$: 1, $S_{p}^{ni}$: 2, $S_{a}$: 2, $S_{m}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3}
[…]
P2 {$S_{g}^{ni}$: 469.16, $S_{c}$: 576.18, $S_{g}$: 651.07, $S_{m}$: 1150.45, $S_{a}$: 1275.04, $S_{p}^{ni}$: 2238.21, $S_{p}$: 3257.35, $S_{r}$: 4274.38, $S_{b}$: 9474.61} {$S_{g}^{ni}$: 1, $S_{c}$: 1, $S_{g}$: 1, $S_{m}$: 2, $S_{a}$: 2, $S_{p}^{ni}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3}
P3 {$S_{g}^{ni}$: 436.33, $S_{g}$: 513.47, $S_{c}$: 650.20, $S_{m}$: 866.61, $S_{r}$: 1305.78, $S_{a}$: 1716.68, $S_{b}$: 1846.56, $S_{p}^{ni}$: 2773.74, $S_{p}$: 4064.03} {$S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{c}$: 1, $S_{m}$: 2, $S_{r}$: 2, $S_{a}$: 2, $S_{b}$: 3, $S_{p}^{ni}$: 3, $S_{p}$: 3}
[…]
P2 {$S_{g}$: 3.72, $S_{c}$: 4.79, $S_{g}^{ni}$: 6.06, $S_{a}$: 11.32, $S_{m}$: 13.36, $S_{p}^{ni}$: 18.94, $S_{p}$: 26.02, $S_{r}$: 42.91, $S_{b}$: 101.20} {$S_{g}$: 1, $S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{a}$: 2, $S_{m}$: 2, $S_{p}^{ni}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3}
P3 {$S_{a}$: 7.87, $S_{g}$: 8.40, $S_{g}^{ni}$: 9.09, $S_{c}$: 9.50, $S_{p}^{ni}$: 16.73, $S_{m}$: 31.36, $S_{r}$: 36.59, $S_{p}$: 37.39, $S_{b}$: 99.72} {$S_{a}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{c}$: 2, $S_{p}^{ni}$: 2, $S_{m}$: 2, $S_{r}$: 3, $S_{p}$: 3, $S_{b}$: 3}
[…]
P2 {$S_{b}$: 0, $S_{g}^{ni}$: 5.31, $S_{c}$: 6.42, $S_{g}$: 8.69, $S_{p}$: 10.81, $S_{m}$: 13.92, $S_{r}$: 17.61, $S_{a}$: 21.39, $S_{p}^{ni}$: 27.63} {$S_{b}$: 1, $S_{g}^{ni}$: 1, $S_{c}$: 1, $S_{g}$: 2, $S_{p}$: 2, $S_{m}$: 2, $S_{r}$: 3, $S_{a}$: 3, $S_{p}^{ni}$: 3}
P3 {$S_{b}$: 0, $S_{g}^{ni}$: 6.48, $S_{g}$: 7.06, $S_{a}$: 7.42, $S_{c}$: 8.80, $S_{m}$: 10.87, $S_{r}$: 16.63, $S_{p}^{ni}$: 35.35, $S_{p}$: 46.50 } {$S_{b}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{a}$: 2, $S_{c}$: 2,$S_{m}$: 2, $S_{r}$: 3, $S_{p}^{ni}$: 3, $S_{p}$: 3 }
[…]
P1 {$S_{c}$: 0.065, $S_{g}^{ni}$: 0.067, $S_{g}$: 0.072, $S_{a}$: 0.084, $S_{m}$: 0.100, $S_{p}$: 0.100, $S_{p}^{ni}$: 0.100, $S_{r}$: 0.830, $S_{b}$: 1.276 } {$S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{a}$: 2, $S_{m}$: 2, $S_{p}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 3, $S_{b}$: 3 }
P2 {$S_{g}$: 0.051, $S_{c}$: 0.053, $S_{g}^{ni}$: 0.060, $S_{a}$: 0.069, $S_{p}^{ni}$: 0.095, $S_{m}$: 0.098, $S_{p}$: 0.100, $S_{r}$: 0.830, $S_{b}$: 1.276 } {$S_{g}$: 1, $S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{a}$: 2, $S_{p}^{ni}$: 2, $S_{m}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
[…]
P1 {$S_{c}$: 5.40, $S_{g}^{ni}$: 5.65, $S_{g}$: 6.13, $S_{p}^{ni}$: 8.87, $S_{p}$: 9.19, $S_{m}$: 9.32, $S_{a}$: 18.36, $S_{r}$: 86.99, $S_{b}$: 947.56 } {$S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{p}^{ni}$: 2, $S_{p}$: 2, $S_{m}$: 2, $S_{a}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
P2 {$S_{g}$: 4.74, $S_{c}$: 4.99, $S_{g}^{ni}$: 5.59, $S_{a}$: 8.24, $S_{p}^{ni}$: 8.49, $S_{m}$: 9.15, $S_{p}$: 9.32, $S_{r}$: 86.87, $S_{b}$: 947.56 } {$S_{g}$: 1, $S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{a}$: 2, $S_{p}^{ni}$: 2, $S_{m}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
[…]
P1 {$S_{m}$: 41.19, $S_{p}$: 44.33, $S_{p}^{ni}$: 46.77, $S_{r}$: 49.62, $S_{g}$: 50.53, $S_{c}$: 52.09, $S_{g}^{ni}$: 53.93, $S_{a}$: 57.08, $S_{b}$: 62.60 } {$S_{m}$: 1, $S_{p}$: 1, $S_{p}^{ni}$: 1, $S_{r}$: 2, $S_{g}$: 2, $S_{c}$: 2, $S_{g}^{ni}$: 3, $S_{a}$: 3, $S_{b}$: 3 }
P2 {$S_{m}$: 41.05, $S_{p}$: 44.02, $S_{p}^{ni}$: 47.67, $S_{r}$: 49.64, $S_{g}$: 49.75, $S_{c}$: 50.79, $S_{g}^{ni}$: 54.43, $S_{a}$: 57.13, $S_{b}$: 62.60 } {$S_{m}$: 1, $S_{p}$: 1, $S_{p}^{ni}$: 1, $S_{r}$: 2, $S_{g}$: 2, $S_{c}$: 2, $S_{g}^{ni}$: 3, $S_{a}$: 3, $S_{b}$: 3 }
Table 2: Final raw scores and ratings based on different metrics computed. Higher ratings indicate higher bias for WRS and PIE %, higher disruption for APE, greater inaccuracy for MASE and SMAPE, and higher accuracy for Sign Accuracy. For simplicity, we denoted the raw scores for accuracy metrics using just the mean value, but standard deviation was also considered for rating. The chosen rating level, L = 3. Overall, across all the settings, system $S_{p}$ exhibited statistical bias in 50 % of cases, confounding bias in 100 % of cases, and disruptive behavior in 50 % of the cases based on APE values.
Metric P $S_{g}$ $S_{g}^{ni}$ $S_{p}$ $S_{p}^{ni}$ $S_{m}$ $S_{c}$ $S_{a}$ $S_{b}$ $S_{r}$
[…]
P1 0.072 $\pm$ 0.123 0.067 $\pm$ 0.178 0.100 $\pm$ 0.125 0.100 $\pm$ 0.143 0.100 $\pm$ 0.076 0.065 $\pm$ 0.189 0.084 $\pm$ 0.282 0.830 $\pm$ 0.639
P2 0.051 $\pm$ 0.047 0.060 $\pm$ 0.085 0.100 $\pm$ 0.088 0.095 $\pm$ 0.097 0.098 $\pm$ 0.074 0.053 $\pm$ 0.092 0.069 $\pm$ 0.217 0.830 $\pm$ 0.639
P3 0.049 $\pm$ 0.045 0.056 $\pm$ 0.052 0.092 $\pm$ 0.074 0.078 $\pm$ 0.078 0.097 $\pm$ 0.071 0.043 $\pm$ 0.048 0.040 $\pm$ 0.037 0.830 $\pm$ 0.640
[…]
P1 6.13 $\pm$ 8.31 5.65 $\pm$ 10.23 9.19 $\pm$ 8.61 8.87 $\pm$ 8.94 9.32 $\pm$ 7.39 5.40 $\pm$ 12.45 18.36 $\pm$ 168.82 86.99 $\pm$ 73.53
P2 4.74 $\pm$ 4.53 5.59 $\pm$ 8.19 9.32 $\pm$ 7.94 8.49 $\pm$ 7.94 9.15 $\pm$ 7.15 4.99 $\pm$ 9.90 8.24 $\pm$ 48.58 86.87 $\pm$ 73.32
P3 4.64 $\pm$ 4.42 5.39 $\pm$ 5.27 8.68 $\pm$ 7.18 7.11 $\pm$ 6.75 9.03 $\pm$ 6.90 4.10 $\pm$ 6.33 3.79 $\pm$ 3.57 86.65 $\pm$ 73.11
Sign Accuracy (%$\uparrow$ P0 52.08 50.41 45.09 47.67 40.70 53.75 60.08 62.60 49.88
P1 50.53 53.93 44.33 46.77 41.19 52.09 57.08 49.62
P2 49.75 54.43 44.02 47.67 41.05 50.79 57.13 49.64
P3 51.34 52.97 44.26 47.50 40.72 51.35 59.98 49.71
Table 3: Performance metrics for test systems across different perturbations. SMAPE and MASE scores are reported as mean $\pm$ standard deviation.

### v3

@@ 4.3. Workflow
5. Experiments and Results
This section introduces the TSFM used in our experiments, baseline models, test data, and evaluation metrics, including two new metrics for perturbations and confounders. We also present the user study design, responses, and findings.
5.1. Experimental Apparatus
5.1.1. Test Models
We evaluate six foundation models (FMs) and three baseline models. Among the FMs, four are used directly, while two (Gemini-V and Phi-3) have both uni-modal and multi-modal variants, making them distinct models. Table 1 provides an overview of the FM architectures. We also consider two ViT-num-spec models, which are vision transformer-based models trained for time-series forecasting task.
Model Mode Size Purpose & Arch. Inf. Time (sec/sample)
Gemini 1.5 Flash Multi 32B GP-1A, 1B, Decoder 1.6 (1A); 10.2 (1B)
Phi-3-vision Multi 4.2B GP-2A, 2B, Enc-Dec 19.7 (1A); 26.6 (1B)
MOMENT-large Uni 385M TS-1, Encoder 0.315
Chronos-T5-small Uni 46M TS-2, Enc-Dec 0.811
ViT-num-spec Multi 86M TS-3, Encoder 6
Table 1. Overview of the architectural details of FMTS. *Best guess in the absence of official information.
Foundation Models:
  1. (1)
    MOMENT ($S_{m}$) (Goswami et al., 2024) is an open-source FM for forecasting, classification, anomaly detection, and imputation in zero-shot and few-shot settings. It is based on the T5-Large encoder (Raffel et al., 2020) and can be fine-tuned if needed.
  2. (2)
    Chronos ($S_{c}$) (Ansari et al., 2024) is a pretrained probabilistic time-series model that tokenizes time-series values using scaling and quantization. It employs a T5 encoder-decoder and is trained via cross-entropy loss. We use Chronos-T5-Small.
  3. (3)
    Gemini-V ($S_{g}$, $S_{g}^{ni}$(Team et al., 2023) is a multi-modal FM designed to process both text and images. We use $S_{g}$ (numeric-only mode) that processes only numerical time-series data, and $S_{g}^{ni}$ (numeric + vision mode) that processes numerical data and time-series line plots.
  4. (4)
    Phi-3 ($S_{p}$, $S_{p}^{ni}$) (Abdin et al., 2024) is a lightweight, state-of-the-art multi-modal FM. We use $S_{p}$ (numeric-only mode) that processes only numerical time-series data, and $S_{p}^{ni}$ (numeric + vision mode) that processes numerical data and time-series line plots. Below is the prompt template used for time-series forecasting with Gemini-V and Phi-3 models (text highlighted in red is omitted for uni-modal forecasting):
    Prompt to Uni-modal and Multi-modal FMs You are a time series forecasting model that only outputs the forecasted numerical values.” ”Input: <time series>” ”Given the input time series for the past 80 time steps and the corresponding time series plot, can you forecast the next 20 time steps? Provide a list of 20 numeric values only. Do not provide any discussion.
    Prompt to Multi-modal FMs for P6 You are a time series forecasting model that only outputs the forecasted numerical values.” ”Input: <time series>” ”Given the input time series from the past 80 time steps and the corresponding sentiment intensity plot, where darker shades indicate more negative sentiment and lighter shades indicate more positive sentiment, Provide a list of 20 numeric values only. Do not provide any discussion.
  5. (5)
    ViT-num-spec Models($S_{v1}$, $S_{v2}$): We employ the ViT-num-spec model (Zeng et al., 2023), which combines a vision transformer with a multimodal time-frequency spectrogram, augmented by the intensities of numeric time series for time series forecasting. This model improves predictive accuracy by leveraging both visual and numerical data. Specifically, it transforms numeric time series into images using a time-frequency spectrogram and utilizes a vision transformer (ViT) encoder with a multilayer perceptron (MLP) head for future predictions.
    Time-Frequency Spectrogram: Building on the method of (Zeng et al., 2023), we use wavelet transforms (Daubechies, 1990) to create time-frequency spectrograms from time series data. Specifically, we employ the Morlet wavelet (Scipy, 2024) with scale $s$ and central frequency $\omega_{0}=5$, as detailed in Equation 7:
    (7) $\psi(x)=\sqrt{\frac{1}{s}}\pi^{-\frac{1}{4}}\exp\left(-\frac{x^{2}}{2s^{2}}\right)\exp\left(j\omega_{0}\frac{x}{s}\right)$
    This method convolves the time series with wavelets at various scales, producing coefficients that indicate signal strength at different frequencies. These magnitudes are visualized in a spectrogram, with higher frequencies at the top and lower ones at the bottom. To retain sign information, a stripe from the standardized numeric time series is added to the top of the spectrogram image. This enhanced image is then used as input for the vision transformer model.
    Vision Transformer: In the next stage, the ViT-num-spec uses a vision transformer with an MLP head for time series forecasting. Input images are segmented into 16 x 16 non-overlapping patches, projected into tokens, and augmented with 1D positional embeddings. The encoder converts these patches into latent representations. For our implementation, 128 x 128 images have price movements in a 16 x 128 top row, with the spectrogram occupying the remaining 112 x 128 space below.
    We trained two variations of the ViT-num-spec model using two real-world datasets and conducted evaluations using a separate dataset. $S_{v1}$ (Pre-COVID training) was trained on S&P 500 stock data from 2000-2014 (46,875 training samples, 46,857 validation samples). $S_{v2}$ (COVID-period training) was trained on data from March 2020–November 2022 (7,478 training samples, 7,475 validation samples).
5.1.2. Baselines
We consider the following baselines:
  1. (1)
    Auto Regressive Integrated Moving Average (ARIMA) ($S_{a}$) is a widely used statistical approach for time series forecasting. It combines three different components: Autoregressive (AR), differencing (I), and moving average (MA) to capture the patterns in the time-series data and predict the next ‘d’ (from section 3) values.
  2. (2)
    Biased system ($S_{b}$) is an extreme baseline biased towards META and GOOG (technology companies), assigning residuals of 0 and 200 respectively, while assigning higher residuals to other companies, representing maximum bias.
  3. (3)
    Random system ($S_{r}$) assigns random price predictions within a company range for contextually meaningful values.
5.1.3. Test Data
5.2. Experimental Evaluation
Figure 4. Plots showing the impact of company and industry as confounders for all the robustness metrics considered. Lower values indicate better robustness.
In this section, we describe the experimental setup used to address the RQs stated in Section 3, the results obtained, and the conclusions drawn from the results. Figure 1(b) shows the causal diagrams used to answer the RQs.
Figure 5. Radar plots showing robustness metrics for all models under different perturbations (P1 through P6) and their average (bottom right). Lower values (i.e., lines closer to the center) indicate lower robustness, while points farther from the center represent better robustness across metrics.
Figure 6. Radar plots showing accuracy metrics for all models under different perturbations (P1 through P6) and their average (bottom right). Lower values (i.e., lines closer to the center) indicate lower accuracy, while points farther from the center represent better accuracy across metrics.
RQ1: Does Sensitive Attribute affect the Residual, even though Sensitive Attribute has no effect on Perturbation?
Setup: In this experiment, the causal link from the Sensitive Attribute to Perturbation is absent, as the perturbation to the stock prices does not depend on the corresponding company name or the industry, i.e., perturbations are applied uniformly across all the data points. We quantify the statistical bias exhibited by the systems by using WRS described in Section 4.2. We perform two different analyses in this experiment: one to measure the discrepancy shown across various industries (WRSIndustry) and another to measure the discrepancy among both the companies (WRSCompany) within the same industry.
Results and conclusion: From Figure 4 and Table 2, most discrepancies can be observed across industries (inter-industry) compared to the discrepancies across companies within each industry. When the input data was subjected to perturbation P2 ($WRS_{avg}$ of 5.55), the systems exhibited more statistical bias. From Figure 5 and Table 2, $S_{a}$ ($WRS_{avg}$ of 3.96) exhibited the least statistical bias, while $S_{p}$ ($WRS_{avg}$ of 6.27) exhibited the highest statistical bias among the systems evaluated under the perturbations considered. Hence, we conclude that Sensitive Attribute affects the Residual, even though Sensitive Attribute has no effect on Perturbation.
Setup: In this experiment, we use PIE % defined in equation 3 to compare the APE (defined in equation 2) before and after deconfounding using the PSM technique as the presence of the confounder opens a backdoor path from Perturbation to Residual through the Confounder. The causal link from Confounder to Perturbation will be valid only if the perturbation applied depends on the value of the confounder (i.e. the company or the industry the specific data points belong to). To ensure the probability of perturbation assignment varies with respect to the Confounder across three distributions (DI1 through DI3) in the case of Industry and six different distributions in the case of Company (DC1 through DC6), we implement weighted sampling. For each distribution, weights are configured so that perturbation groups P1 through P6 have a twofold higher likelihood of selection compared to P0 for specific values of the confounder. For example, META in DC1, GOOG in DC2, and so on. This strategy highlights significant cases, although other combinations are possible for further exploration.
Results and Conclusion: Figure 4 shows that selecting Industry as the confounder leads to greater confounding bias in the systems. In Figure 5, $S_{g}^{ni}$ ($PIE_{avg}\%$ of 1107.66) exhibited the least confounding bias, while $S_{p}^{ni}$ ($PIE_{avg}\%$ of 2778.06) exhibited the most. Systems showed more confounding bias under perturbation P5 ($PIE_{avg}\%$ of 3646.20). Therefore, the Confounder affects the relationship between Perturbation and Residual, particularly when the Confounder influences the Perturbation.
RQ3: Does Perturbation affect the Residual when Sensitive Attribute may have an effect on Residual?
Setup: The experimental setup in this experiment is the same as that for answering RQ2. To compute the APE, we used PSM described in Section 4.2. PSM allows us to effectively determine the effect of Perturbation on the Residual. For instance, if two matched points belong to the same company but only one was perturbed, any difference in their residuals can be directly attributed to the perturbation itself rather than to other confounding factors. This method provides a clear understanding of the true impact of the Perturbation on the Residual. As our rating method aims to bring out the worst possible behavior of the systems, we take the MAX(APE) as the raw score that is used to compute the final ratings.
Results and Conclusion: It is undesirable to have a higher APE, as it implies that the perturbation applied can have a significant impact on the residuals of different systems. From Figure 4, when Industry was considered as the confounder, it led to a higher APE. As the outcome of $S_{b}$ depended on the Company (and varied from one company to another), the perturbation did not have any effect on the system. Whereas, when Industry was considered as the confounder, the perturbation appeared influential, resulting in a high APE for $S_{b}$. From Figure 5 and Table 2, perturbations had the least impact on $S_{g}^{ni}$ ($APE_{avg}$ of 5.89) and highest impact on $S_{a}$ ($APE_{avg}$ of 27.73). Among all the perturbations, P1 ($APE_{avg}$ of 20.25) was the most disruptive. Hence, Perturbation affects the Residual when Sensitive Attribute may have an effect on Residual.
RQ4: Does Perturbations degrade the accuracy of S?
Setup: In this experiment, we compute the three accuracy metrics widely used in for the task of financial time-series forecasting (Makridakis et al., 2022), which were summarized in Section 4.2.
Results and Conclusion: From Figure 6, $S_{c}$ exhibited the highest amount of forecasting accuracy in terms of SMAPE (average of 0.05) and MASE (average of 4.67), while $S_{a}$ outperformed all other systems in terms of sign accuracy (average of 58.57). $S_{b}$ consistently predicted the correct directional movement of stock prices, exhibiting high sign accuracy as it was designed to adjust residuals based on specific company stock prices. Perturbation P6 caused the highest decline in SMAPE (average of 0.39) and MASE (average of 176.43), while P1 caused the highest decline in sign accuracy (average of 49.86). Hence, Perturbations degrade the accuracy of S.
5.3. Overall Performance Comparison